# Can the 3 neutrino masses really be found using SN 1987A data?

### v6

@@ II Finding $m^{2}$ of SN 1987A neutrinos
Obviously steeper sloped lines yield smaller masses, a photon lies on a vertical line, and $m^{2}<0$ neutrinos lie on a line through the origin having negative slope.
@@ III Fitting 37 data points to 3 lines
Thirty-seven neutrinos is not a very large number, but the distinctiveness of a pattern counts for far more than the sheer number of events. Thus, the $\Omega^{-}$ particle which led to quarks was confirmed in 1964 at Brookhaven National Laboratory by observing just one single“Gold Standard” event. (Barnes1964, ) In the present case, the pattern found when the SN 1987A data is plotted (see Fig.1) is also very distinctive since every one of the neutrinos from SN 1987A lies on or next to one of three straight lines through the origin as predicted, and only one data point very near the origin is within an error bar width of two lines. The least squares fit of 37 data points to three lines through the origin has three free parameters, $S_{j}$ the slopes of the three straight lines in Fig. 1 yielding 34 degrees of freedom. The parameter H, the size of the constant horizontal error bar on each data point in Fig.1 is also adjusted. Since each data point has both x and y error bars, we can then convert the x error into an ”effective” y error by multiplying the x uncertainty by the slope of the line, and combine it with the y uncertainty in quadrature. If a data point lies close to two straight lines, we simply choose the line that gives the smaller contribution to chi square.
@@ III.1 The best fit
Table 1: Seven “extra” Kamiokande II data points extracted from Fig. 4f of Hirata et al. (Hirata1988, ). Thse seven data points have $N_{hits}>24$ and $13<t<25$ seconds. The energy for each data point was estimated based on: $E=0.36N_{hits}.$ The uncertainty in the energy (before 1988) was $\Delta E/E=22\%/\sqrt{E/10MeV}$ (Hirata1989, )
III.2 The 7 extra Kamiokande II neutrinos
One problem in fitting the SN 1987A data to three straight lines is deciding what is the maximum time to use in the fit. This is important since after many seconds neutrinos are increasingly likely to be just background and not be associated with SN 1987A. Ref.  (Hirata1988, ) for Kamiokande II arbitrarily chose $t=12$ sec as the time cutoff. This choice resulting in their tabulating data for 12 SN 1987A neutrinos. It was concerning to the author, however, that the last data point at $t=12$ seconds in Fig. 1 is $1.9\sigma$ above the fitted line. This concern motivated the desire to add more consecutive data points beyond $t=12$ sec, and seven of these were extracted by the author from ref. (Hirata1988, ) as explained in Table I. These data points are very unlikely to be due to background which lies below 20 hits, i.e. 20 photomultipliers activated, and adding them has had little effect on the goodness of the fit to three straight lines or their slopes.
@@ III.3 An omission in most sterile neutrino searches
The existence of sterile neutrinos is highly controversial, and there are probably more negative search results than positive ones. Thus, neither MicroBooNE, nor IceCube, nor KATRIN nor DayaBay report any evidence for them. (Micro2025, ; Aartsen2020, ; KATRIN2025, ; An2024, ). Nevertheless, an omission in these and other experiments is that most have not sought evidence for the presence of multiple sterile neutrinos simultaneously. There are several possible remedies for this omission, the simplest being to search for two or three sterile neutrinos in the same fit to the data, and another being to find a place where nature causes a spatial separation between light and heavy sterile neutrinos.
To date the only search for the pair of $m^{2}>0$ sterile neutrinos suggested by the SN 1987A data has been carried out by M. H. Chan and the author. (Chan2014, ) Essentially, we first derived an equation of state for the sterile neutrinos dependent on their mass, and then separately found the dark matter radial profile in both the Milky Way and four groups of galaxies. Finally, the mass was varied to find the value that best fit the observed profile in each case. The results of these fits showed that the two $m^{2}>0$ values from the 3-mass SN 1987A fit would indeed fit the galaxy clusters (lighter mass) and the Milky Way (heavier mass). This result offers a confirmation of the $m_{1}=21.6eV/c^{2}$ and $m_{2}=2.70eV/c^{2}$ masses deduced from the SN 1987A three mass fit.
@@ IV.2 Why are the LSD data so bunched in time?
If the LSD burst of 5 neutrinos is genuine, it may look strange in Fig. 1 that while the other data points lie all along their respective lines the five LSD neutrinos are all bunched together in time at t=484t=-484 min or -17 ksec. The three straight line result would be more believable perhaps if the five neutrinos yielded data points at five widely spaced points on the negatively sloped line. But recall that the $t<0$ data have a kilosecond time scale. If the five LSD data points were spread out along the negative sloped line and not bunched in time they never could have been distinguished from background because researchers were looking for a concentration of neutrino events in a seconds-long time interval, i.e., a burst.
@@ IV.5 Best size of the horizontal error bars?
The main way the good fit to three straight lines could be a mirage is that the neutrinos were not emitted simultaneously as our derivation requires, meaning that the size of the emissions time window was too large for the method to work. If all neutrinos were emitted simultaneously (a zero width time window) the recorded time of each neutrino would accurately represent its emission time, and there would be no need for having any horizontal error bars. If the time window were an amount $\Delta t$ then each neutrino might have its arrival time t randomly shifted to the right by a time up to that amount, and a horizonal error bar $H\approx\Delta t$ would give a good fit to three straight lines. Thus, we can find the size of the emission time window by seeing for what horizontal error bar width H, the goodness of the fit probability significantly improves. Clearly from Fig. 2 it is about 0.5 to 1.0 seconds.
IV.6 The time window for neutrino emissions?
We can estimate the actual size of the time window for neutrino emissions using a simulation. In 2025 Bozza et al. have calculated the neutrino luminosity from SN 1987A using a new model they recently developed. (Bozza2025, ) They claim the analysis is in many ways more refined and accurate than previous ones with respect to the temporal structure of the neutrino luminosity. The functional form of their computed luminosity has 9 parameters that are determined by a least squares fit to the SN 1987A data, and their result is shown in Fig. 4. Note that the maximum neutrino emission occurs at about $T=100ms=0.1sec$ and the bulk of the emissions occur for $t<1$ second. Their best fit parameter values are all in accordance with current supernova simulations, and they find surprisingly that $95\%$ of the analyzed events (even those at $t>>1$ sec) are signal rather than background. These results all buttress our case that the SN 1987A neutrinos were mostly emitted during a 0.5 to 1.0 sec time window, and that our use of $\pm 0.5$ sec for the optimal horizontal uncertainty in our least squares fit to three straight lines was justified. A 0.5 to 1.0 sec time window is apparently small enough not to badly degrade the fit of 37 neutrino data points to three straight lines.
@@ IV.12 Multiple sterile neutrinos in oscillation data?
The need for doing searches for multiple sterile neutrinos in a single search has already been noted in section III.3. How would one go about seeking evidence for both $m_{1}^{2}=2.70^{2}eV^{2}/c^{4},$ and $m_{2}=21.6eV^{2}/c^{4}$ in oscillation data, bearing in mind that the oscillation wavelength for $m_{2}$ would be $21.6^{2}/2.7^{2}=64$ times shorter than the $m_{1}$ sterile neutrinos? It is unlikely that such short wavelength oscillations are detectable in the data of Neutrino 4, but if they were, one approach to find them would involve setting up a grid of values for the two masses $m_{1},m_{2}$ and in each case use the data to find the likelihood probability of the fit to the data for that pair of mass values. One could then see if there is a peak right at a small region of the grid near $(m^{2}_{1},m_{2}^{2})=(2.70^{2},21.6^{2})eV^{2}/c^{2}.$ A more likely possibility would be the IceCube experiment. IceCube has sought evidence for $\nu_{\mu}$ and $\bar{\nu_{\mu}}$} disappearance oscillations in 10.7 years of data. (Abbasi2025, ) Even though IceCube reported a negative result their best fit they did show a hint of a value consistent with the $m_{1}=2.70eV/c^{2}$ mass which was strongly preferred over the no oscillation solution by 40 to one $p=2.5\%.$

### v7

@@ II Finding $m^{2}$ of SN 1987A neutrinos
Obviously steeper sloped lines yield smaller masses, a photon lies on a vertical line, and $m^{2}<0$ neutrinos lie on a line through the origin lying in the $t<0$ region.
@@ III Fitting 37 data points to 3 lines
Thirty-seven neutrinos is not a very large number, but the distinctiveness of a pattern counts for far more than the sheer number of events. Thus, the $\Omega^{-}$ particle which led to quarks was confirmed in 1964 at Brookhaven National Laboratory by observing just one single“Gold Standard” event. (Barnes1964, ) In the present case, the pattern found when the SN 1987A data is plotted (see Fig.1) is also very distinctive since every one of the neutrinos from SN 1987A lies on or next to one of three straight lines through the origin as predicted, and only one data point very near the origin is within an error bar width of two lines. The least squares fit of 37 data points to three lines having a common $t=0$ origin has three free parameters, $S_{j}$ the slopes of the three straight lines in Fig. 1 yielding 34 degrees of freedom. The parameter H, the size of the constant horizontal error bar on each data point in Fig.1 is also adjusted. Since each data point has both x and y error bars, we can then convert the x error into an ”effective” y error by multiplying the x uncertainty by the slope of the line, and combine it with the y uncertainty in quadrature. If a data point lies close to two straight lines, we simply choose the line that gives the smaller contribution to chi square.
@@ III.1 The best fit
Table 1: Seven “extra” Kamiokande II data points extracted from Fig. 4f of Hirata et al. (Hirata1988, ). These seven data points have $N_{hits}>24$ and $13<t<25$ seconds. The energy for each data point was obtained by the author based on the linear relation: $E=0.363N_{hits}$ to an accuracy of $5\%$ (Ehrlich2018, ). The uncertainty in the energy (before 1988) was $\Delta E/E=22\%/\sqrt{E/10MeV}$ (Hirata1989, )
III.2 The 7 “extra” Kamiokande II neutrinos
One problem in fitting the SN 1987A data to three straight lines is deciding what is the maximum time to use in the fit. This is important since after many seconds neutrinos are increasingly likely to be just background and not be associated with SN 1987A. Ref.  (Hirata1988, ) for Kamiokande II arbitrarily chose $t=12$ sec as the time cutoff. This choice resulting in their tabulating data only for 12 SN 1987A neutrinos. It was concerning to the author, however, that the last data point at $t=12$ seconds in Fig. 1 is $1.9\sigma$ above the fitted line. This concern motivated the desire to add more consecutive data points beyond $t=12$ sec, and seven of these were extracted by the author from ref. (Hirata1988, ) as explained in Table I. These data points are very unlikely to be due to background which lies below 20 hits, i.e. 20 photomultipliers activated, and adding them improved the goodness of the fit to three straight lines.
@@ III.3 An omission in most sterile neutrino searches
The existence of sterile neutrinos is highly controversial, and there are probably more negative search results than positive ones. Thus, neither MicroBooNE, nor IceCube, nor KATRIN nor DayaBay report any evidence for them. (Micro2025, ; Aartsen2020, ; KATRIN2025, ; An2024, ). Nevertheless, an omission in these and other experiments is that most have not sought evidence for the presence of multiple sterile neutrinos simultaneously, i.e., they do $3+1$ and not $3+2$ fits to data. There are two possible remedies for this omission, the simplest being to search for two or three sterile neutrinos in the same fit, and another one being to find a place where nature causes a spatial separation between light and heavy sterile neutrinos.
To date the only search for the pair of $m^{2}>0$ sterile neutrinos suggested by the SN 1987A data has been carried out by M. H. Chan and the author. (Chan2014, ) Sterile neutrinos are a natural possibility for the identity of dark matter in the galaxy and clusters of galaxies. Essentially, the authors first derived an equation of state for the sterile neutrinos dependent on their mass, and then separately found the dark matter radial profile in both the Milky Way and four groups of galaxies. Finally, the mass was varied to find the value that best fit the observed profile in each case. The results of these fits showed that the two $m^{2}>0$ values from the 3-mass SN 1987A fit would indeed fit the galaxy clusters (lighter mass) and the Milky Way (heavier mass). This result offers a confirmation of the $m_{1}=21.6eV/c^{2}$ and $m_{2}=2.70eV/c^{2}$ masses deduced from the SN 1987A three mass fit.
@@ IV.2 Why are the LSD data so bunched in time?
If the LSD burst of 5 neutrinos is genuine, it may look strange in Fig. 1 that while the other data points lie all along their respective lines the five LSD neutrinos are all bunched together in time at t=484t=-484 min or -17 ksec. The three straight line result would be more believable perhaps if the five neutrinos yielded data points at five widely spaced points on the negatively sloped line. But recall that the $t<0$ data have a kilosecond time scale. If the five LSD data points were spread out along the LSD line and not bunched in time they never could have been distinguished from background because researchers were looking for a concentration of neutrino events in a seconds-long time interval, i.e., a burst. The fact that the 5 LSD neutrinos were recorded at -17 ksec does, of course, not mean they went backward in time, only that they travelled faster than light, assuming they started at $t=0$ along with all the other three neutrino bursts. Because of that bunching at the left end of the LSD line the uncertainty in the slope (and the value of $m^{2}$) is $18\%$ or $m^{2}=-192,000\pm 34,000eV^{2}/c^{4}$.
@@ IV.5 Best size of the horizontal error bars?
The main way the good fit to three straight lines could be a mirage is that the neutrinos were not emitted simultaneously as our derivation requires, meaning that the size of the emissions time window was too large for the method to work. If all neutrinos were emitted simultaneously (a zero width time window) the recorded time of each neutrino would accurately represent its emission time, and there would be no need for having any horizontal error bars. If the time window were an amount $\Delta t$ then each neutrino might have its arrival time t randomly shifted to the right by a time up to that amount, and a horizonal error bar $H\approx\Delta t$ would give a good fit to three straight lines. Thus, we can find the size of the emission time window by seeing for what horizontal error bar width H, the goodness of the fit probability significantly improves. Clearly from Fig. 2 it is about 0.5 to 1.0 seconds. Note that the time any individual neutrino is recorded is, of course, known to much greater precision than 0.5 seconds. However, the meaning of H takes into account the uncertainty of when the burst of neutrinos this neutrino was part of occurred, which is primarily determined by the choice of $t=0$ for the first arriving beutrino in a detector.
IV.6 A simulation of SN 1987A neutrino emissions
In 2025 Bozza et al. have calculated the neutrino luminosity from SN 1987A using a new model they recently developed. (Bozza2025, ) They claim the analysis is in many ways more refined and accurate than previous ones with respect to the temporal structure of the neutrino luminosity. The functional form of their computed luminosity has 9 parameters that are determined by a least squares fit to the SN 1987A data, and their result is shown in Fig. 4. Note that the maximum neutrino emission occurs at about $T=100ms=0.1sec$ and the bulk of the emissions occur for $t<1$ second. Their best fit parameter values are all in accordance with current supernova simulations, and they find surprisingly that $95\%$ of the analyzed events (even those at $t>>1$ sec) are signal rather than background.
These Bozza et al. results buttress our case that the SN 1987A neutrinos were mostly emitted during a 0.5 to 1.0 sec time window, and that our use of $\pm 0.5$ sec for the optimal horizontal uncertainty in our least squares fit to three straight lines was justified. A 0.5 to 1.0 sec time window is apparently small enough not to badly degrade the fit of 37 neutrino data points to three straight lines. However, the use of as many as 9 free parameters in the Bozza spectral fit is a serious drawback, since with enough free parameters one can fit virtually any shape. In contrast, our fit to the $m^{2}>0$ neutrino spectrum used only two parameters, i.e., the two $m^{2}>0$ masses. Moreover, as long as the spread in neutrino emission times is small compared to the horizontal error bars used (as implied by Fig. 2), relativistic kinematics requires that neutrinos of a given mass lie on straight lines having a specific slope, and there is no similar reqirement in the Bozza analysis.
@@ IV.12 Multiple sterile neutrinos in oscillation data?
The need for doing searches for multiple sterile neutrinos in a single search has already been noted in section III.3. How would one go about seeking evidence for both $m_{1}^{2}=2.70^{2}eV^{2}/c^{4},$ and $m_{2}=21.6eV^{2}/c^{4}$ in oscillation data, bearing in mind that the oscillation wavelength for $m_{2}$ would be $21.6^{2}/2.7^{2}=64$ times shorter than the $m_{1}$ sterile neutrinos? It is unlikely that such short wavelength oscillations are detectable in the data of Neutrino 4, but if they were, one approach to find them would involve setting up a grid of values for the two masses $m_{1},m_{2}$ and in each case use the data to find the likelihood probability of the fit to the data for that pair of mass values. One could then see if there is a peak right at a small region of the grid near $(m^{2}_{1},m_{2}^{2})=(2.70^{2},21.6^{2})eV^{2}/c^{2}.$ A more likely possibility would be the IceCube experiment. IceCube has sought evidence for $\nu_{\mu}$ and $\bar{\nu_{\mu}}$} disappearance oscillations in 10.7 years of data. (Abbasi2025, ) Even though IceCube reported a negative result their best fit they did show a hint of a value consistent with the $m_{1}=2.70eV/c^{2}$ mass which was strongly preferred over the no oscillation solution by 40 to one $p=2.5\%.$ To date, however, IceCube has only done a limited amount of searching for multiple sterile neutrinos. (Cabrera2024, )